# FindingDory: A Benchmark to Evaluate Memory in Embodied Agents

### v1

@@ D.4 Text Agent
We implement the the text-based VLM agent using the Qwen2.5-VL [5] model. For every chunk_size number of frames in the video, we generate a description using the VLM. We then attach the frame indices and time of day information to the generated summary for the frame window in consideration. We use the prompt shown in Fig. 9 to generate the structured JSON outputs for each chunk in the video. We then concatenate all the generated JSON outputs and pass it to the Qwen model to perform text-only reasoning over the entire text-based history of the video. The text-based reasoning prompt is provided in Fig. 10. We note that using the same Qwen model for generating the individual text summaries and performing the final text-based reasoning helps in optimizing inference load as we maintain only a single instance of the VLM through the entire evaluation.
Text Agent(VLM) Task Prompt You are an expert and intelligent question answering agent. You will be shown a stream of {chunk_size} images taken by a robot while navigating around a simulated house and picking and placing objects. Your job is to describe the image based on the output format. Make sure your response is all within the JSON format.
Output Format (should be in JSON format):
@@
  "picking_placing_or_navigating": str("picking", "placing" or "navigating").
@@
Where:
- Room name: The name of the room the robot is in.
- Picking, placing, or navigating: Whether the robot is picking or placing an object, or navigating around the house.
- Object being manipulated: The object that the robot is picking or placing (if relevant).
- Receptacle being manipulated: The receptacle that the robot is picking up from or placing an object into (if relevant).
- Other objects in scene: Objects present in the scene besides the main object and the receptacle.
Note:
- The robot uses a magic grasp action to pick up an object, where a gripper goes close to the object and the object gets magically picked up.
Figure 9: System Prompt for Qwen2.5-VL model to generate text summaries based on chunk_size frames
Text Agent(LLM) Task Prompt You are an expert and intelligent question answering agent. You will be provided a list of textual descriptions of the environment created by an agent while navigating around a house and picking and placing objects. The descriptions correspond to logs made after collecting and viewing a small chunk of frames, and contain information about which frame numbers they correspond to. You will also be provided a goal that now needs to be accomplished, for which you will need to use the history to decide where to go. Your job is to identify the desired frame index to navigate to, based on the provided task. Make sure your response is all within the JSON.
Agent’s History information:
- Room name: The name of the room the agent was in.
- Picking, placing, or navigating: Whether the robot was picking or placing an object, or navigating around the house.
- Object being manipulated: The object that the robot was picking or placing (if relevant).
- Receptacle being manipulated: The receptacle that the robot was picking up from or placing an object into (if relevant).
- Other objects in scene: Objects present in the scene besides the main object and the receptacle.
Output Format (should be in JSON format):
@@
Where:
- Chain of Thought: A detailed explanation of your thought process in determining the frame indices.
- Frame Indices: A list of one or more frame indices relevant to accomplishing the goal.
- If its a single object/place task, output only one frame index.
- If task requires the agent to go to multiple places, output one frame index per object/place. (Do not use ellipsis)
Agent’s history: history
The goal is {goal}
Figure 10: System Prompt for Qwen2.5-VL model to perform textual reasoning over the concatenated textual summaries
@@ D.5 Supervised Finetuning Baseline
We leverage the FindingDory training split which consists of $82174$ unique task instantiations across the categories mentioned in Table 2. Each video-instruction pair is coupled with a list of ground-truth frame indices that solve the task. For multi-goal tasks, each subgoal has a corresponding ground-truth frame indices sublist. Since the FindingDory video sequences can be extremely long, we uniformly subsample all videos to $96$ frames. The ground truth frame indices are also subsampled to match the shorter video lengths to ensure the model chooses from within the subsampled frames.
Training Details. We conduct full-finetuning of the Qwen/Qwen2.5-VL-3B model using the open-source huggingface/trl library. For hyperparameter optimization, we utilized a smaller representative split of the training dataset with $36274$ samples and conducted a grid search over the learning rate, num_epochs and weight_decay parameters. The best checkpoints from all runs were selected by evaluating performance on all video-instruction pairs in the full validation split. This “offline” evaluation is conducted without instantiating the simulator by directly comparing the model’s predicted keyframe indices with the ground-truth keyframe indices using exact string matching (thus, we dont compute metrics proposed in Section 4.2). We use the cosine learning rate decay scheduler for all experiments. For the final training on the larger dataset, we found that a learning rate of $5\times 10^{-6}$ over $5$ epochs with 00 weight decay worked best. For all experiments, we train in bfloat16 format using 8 A40 GPUs on a single node with a batch size of $1$ and gradient_accumulation of $4$ providing an effective batch_size of $32$ samples. To further optimize training, we use gradient_checkpointing and highly optimized flash_attention implementation. In this setup, the full training run takes $\approx 120$ hrs to complete.
@@ E.1 Image-Goal Navigation Agent
Policy Training. We train the image-goal navigation policy using the architecture and training recipe described in [46]. Specifically, we employ an end-to-end reinforcement learning policy trained using DDPPO, which predicts discrete navigation actions conditioned on visual RGB inputs and a goal image. The agent utilizes a ViT based visual encoder along with a 2-layer LSTM backbone. Similar to OVRL-v2 [46], the output patch representations from the ViT are reshaped into a 3D grid and downsampled to a lower dimension using a convolutional layer called the compression layer. However, we slightly deviate from OVRL-v2 by first concatenating the patch representations from the current and goal images before passing them through the compression layer.
The policy training occurs in the training scene split of the HSSD dataset [15], consisting of $1166$ episodes distributed across $111$ scenes (a subset of the original $125$ training scenes). In the Habitat simulator, the agent is modeled as a Hello Robot Stretch [14] with a height of $1.41$ m and a cylindrical base radius of $0.25$ m. The agent’s RGB sensor is positioned at a height of $1.31$ m, with a resolution of $160\times 120$ pixels and a horizontal field of view of $43^{\circ}$. During training, each episode has a maximum step budget of $1000$ steps, with success defined by invoking the STOP action within $1.0$ m of the goal image position.
As our visual encoder, we use the VC-1-Base [22] model which was previously finetuned with ImageNav on a smaller training scene split from the HSSD dataset. Empirically, we found that freezing the visual encoder after this targeted finetuning significantly accelerates the overall training performance. During the VC-1 finetuning, we use a smaller learning rate of $1.5\times 10^{-6}$ for the encoder.
We train the agents for a total of $500$M timesteps on 32 A40 GPUs running 32 parallel environments each (1024 envs in total). Following the approach from [22], we collect $64$ steps of experience and subsequently perform 2 PPO epochs with a mini-batch size of $2$. The reward function employed is the improved formulation proposed by [46], using parameters: success weighting $c_{s}=5.0$, angle success weighting $c_{a}=5.0$, goal radius $r_{g}=1.0$, angle threshold $\theta_{g}=25^{\circ}$, and slack penalty $\gamma=-0.002$. We optimize using AdamW with a learning rate of $2.5\times 10^{-4}$ and weight decay $10^{-6}$.
@@ E.2 Explicit Mapping Agent
Planner. When the high-level goal agent selects a particular goal frame, we lookup the corresponding target 2D pose in the global map. Simlar to the local navigation policy employed in [8], we use the Fast Marching Method [34] to generate the shortest path to the goal pose. We use one of four discrete navigation actions (see Section 3.3) to reach each waypoint on the generated shortest path.
Appendix F Qualitative Examples
@@
In this section, we include representative qualitative examples from the FindingDory task suite. In Figure 12, we provide a birds-eye view depiction of the trajectory executed during experience collection (see Section B.2) in a validation episode involving 5 pick-place interaction routines. For this episode, the collected video consists of $1308$ frames in total. We provide detailed responses (and meta-analysis) from various baselines across the various task categories in Figures 14, 15, 16, 17, 18 and 19. In the following, we briefly discuss the different failure modes encountered in the generated responses across the 3 high-level task categories defined in Table 2.
@@
Single-Goal Spatial Tasks. We observe that for the Object Recall task (Fig. 14), the models are able to identify a frame with the target chest of drawers but do not select the closest frames to ensure success. In the Interaction based tasks (Figs. 14 and 15), we observe models fail to accurately list the sequence of interaction events which can be partially attributed to supplying only $96$ subsampled frames from the original video. In one case, GPT-4o instead focuses on a “distractor” target object (Fig. 15). All models perform poorly on the Spatial Relationship tasks (Fig. 15) as they are unable to coherently reason about the spatial layout of the house in the provided video.
@@
Figure 14: FindingDory Spatial Tasks for the sample episode in Figure 12. Each VLM model uses $96$ subsampled frames. Text Agent processes full video as defined in Sec. D.4.
Figure 15: FindingDory Spatial Tasks for the sample episode in Figure 12. Each VLM model uses $96$ subsampled frames. Text Agent processes full video as defined in Sec. D.4.
Figure 16: FindingDory Spatial Tasks for the sample episode in Figure 12. Each VLM model uses $96$ subsampled frames. Text Agent processes full video as defined in Sec. D.4.
Figure 17: FindingDory Temporal Tasks for the sample episode in Figure 12. Each VLM model uses $96$ subsampled frames. Text Agent processes full video as defined in Sec. D.4.
Figure 18: FindingDory Temporal Tasks for the sample episode in Figure 12. Each VLM model uses $96$ subsampled frames. Text Agent processes full video as defined in Sec. D.4.
Figure 19: FindingDory Multi-Goal Tasks for the sample episode in Figure 12. Each VLM model uses $96$ subsampled frames. Text Agent processes full video as defined in Sec. D.4.

### v2

@@ D.4 Text Agent
We implement the the text-based VLM agent using the Qwen2.5-VL (Bai et al., 2025) model. For every chunk_size frames in the video, we generate a description using the VLM. We then attach the frame indices and time of day information to the generated summary for the frame window in consideration. We use the prompt shown in Fig. 13 to generate the structured JSON outputs for each chunk in the video. We then concatenate all the generated JSON outputs and pass it to the Qwen model to perform text-only reasoning over the entire text-based history of the video. The text-based reasoning prompt is provided in Fig. 14. We note that using the same Qwen model for generating the individual text summaries and performing the final text-based reasoning helps in optimizing inference load as we maintain only a single instance of the VLM through the entire evaluation.
Text Agent(VLM) Task Prompt You are an expert and intelligent question answering agent. You will be shown a stream of {chunk_size} images taken by a robot while navigating around a simulated house and picking and placing objects. Your job is to describe the image based on the output format. Make sure your response is all within the JSON format. Output Format (should be in JSON format):
@@
  "picking_placing_or_navigating": str("picking", "placing" or "nav").
@@
Where: - Room name: The name of the room the robot is in. - Picking, placing, or navigating: Whether the robot is picking or placing an object, or navigating around the house. - Object being manipulated: The object that the robot is picking or placing (if relevant). - Receptacle being manipulated: The receptacle that the robot is picking up from or placing an object into (if relevant). - Other objects in scene: Objects present in the scene besides the main object and the receptacle. Note: - The robot uses a magic grasp action to pick up an object, where a gripper goes close to the object and the object gets magically picked up.
Figure 13: Prompt to generate text summaries for chunk_size frames with Qwen2.5-VL
Text Agent(LLM) Task Prompt You are an expert and intelligent question answering agent. You will be provided a list of textual descriptions of the environment created by an agent while navigating around a house and picking and placing objects. The descriptions correspond to logs made after collecting and viewing a small chunk of frames, and contain information about which frame numbers they correspond to. You will also be provided a goal that now needs to be accomplished, for which you will need to use the history to decide where to go. Your job is to identify the desired frame index to navigate to, based on the provided task. Make sure your response is all within the JSON. Agent’s History information: - Room name: The name of the room the agent was in. - Picking, placing, or navigating: Whether the robot was picking or placing an object, or navigating around the house. - Object being manipulated: The object that the robot was picking or placing (if relevant). - Receptacle being manipulated: The receptacle that the robot was picking up from or placing an object into (if relevant). - Other objects in scene: Objects present in the scene besides the main object and the receptacle. Output Format (should be in JSON format):
@@
Where: - Chain of Thought: A detailed explanation of your thought process in determining the frame indices. - Frame Indices: A list of one or more frame indices relevant to accomplishing the goal. - If its a single object/place task, output only one frame index. - If task requires the agent to go to multiple places, output one frame index per object/place. (Do not use ellipsis) Agent’s history: {history} The goal is {goal}
Figure 14: Prompt to perform reasoning over the combined textual summaries with Qwen2.5-VL
@@ D.5 Supervised Finetuning Baseline
We leverage the FindingDory training split which consists of $82174$ unique tasks across the categories mentioned in Table 2. Each video-instruction pair is coupled with a list of ground-truth frame indices that solve the task. For multi-goal tasks, each subgoal has a corresponding ground-truth frame indices sublist. Since the FindingDory video sequences can be extremely long, we uniformly subsample all videos to $96$ frames. The ground truth frame indices are also subsampled to match the shorter video lengths to ensure the model chooses from within the subsampled frames.
Training Details. We conduct full-finetuning of the Qwen/Qwen2.5-VL-3B model using the open-source huggingface/trl library. For hyperparameter optimization, we utilized a smaller representative split of the training dataset with $36274$ samples and conducted a grid search over the learning rate, num_epochs and weight_decay parameters. The best checkpoints from all runs were selected by evaluating performance on all video-instruction pairs in the full validation split. This “offline” evaluation is conducted without instantiating the simulator by directly comparing the model’s predicted keyframe indices with the ground-truth keyframe indices using exact string matching (thus, we dont compute metrics proposed in Section 4.2). We use the cosine learning rate decay scheduler for all experiments. For the final training on the larger dataset, we found that a learning rate of $5\times 10^{-6}$ over $5$ epochs with 0 weight decay worked best. For all experiments, we train in bfloat16 format using 8 A40 GPUs on a single node with a batch size of 11 and gradient_accumulation of $4$ - providing an effective batch_size of $32$ samples. To further optimize training, we use gradient_checkpointing and highly optimized flash_attention implementation. In this setup, the full training run takes $\approx 120$ hrs to complete.
@@ E.1 Image-Goal Navigation Agent
Policy Training. We train the image-goal navigation policy using the architecture and training recipe described in  Yadav et al. (2023). Specifically, we employ an end-to-end reinforcement learning policy trained using DDPPO, which predicts discrete navigation actions conditioned on visual RGB inputs and a goal image. The agent utilizes a ViT based visual encoder along with a 2-layer LSTM backbone. Similar to OVRL-v2 (Yadav et al., 2023), the output patch representations from the ViT are reshaped into a 3D grid and downsampled to a lower dimension using a convolutional layer called the compression layer. However, we slightly deviate from OVRL-v2 by first concatenating the patch representations from the current and goal images before passing them through the compression layer.
The policy training occurs in the training scene split of the HSSD dataset (Khanna* et al., 2023), consisting of $1166$ episodes distributed across $111$ scenes (a subset of the original $125$ training scenes). In the Habitat simulator, the agent is modeled as a Hello Robot Stretch (Kemp et al., 2022) with a height of $1.41$ m and a cylindrical base radius of $0.25$ m. The agent’s RGB sensor is positioned at a height of $1.31$ m, with a resolution of $160\times 120$ pixels and a horizontal field of view of $43^{\circ}$. During training, each episode has a maximum step budget of $1000$ steps, with success defined by invoking the STOP action within $1.0$ m of the goal image position.
As our visual encoder, we use the VC-1-Base (Majumdar et al., 2023) model which was previously finetuned with ImageNav on a smaller training scene split from the HSSD dataset. Empirically, we found that freezing the visual encoder after this targeted finetuning significantly accelerates the overall training performance. During the VC-1 finetuning, we use a smaller learning rate of $1.5\times 10^{-6}$ for the encoder.
We train the agents for a total of 500500M timesteps on 32 A40 GPUs running 32 parallel environments each (1024 envs in total). Following the approach from  (Majumdar et al., 2023), we collect $64$ steps of experience and subsequently perform 2 PPO epochs with a mini-batch size of $2$. The reward function employed is the improved formulation proposed by Yadav et al. (2023), using parameters: success weighting $c_{s}=5.0$, angle success weighting $c_{a}=5.0$, goal radius $r_{g}=1.0$, angle threshold $\theta_{g}=25^{\circ}$, and slack penalty $\gamma=-0.002$. We optimize using AdamW with a learning rate of $2.5\times 10^{-4}$ and weight decay $10^{-6}$.
@@ E.2 Explicit Mapping Agent
Planner. When the high-level goal agent selects a particular goal frame, we lookup the corresponding target 2D pose in the global map. Similar to the local navigation policy employed in Chang et al. (2023), we use the Fast Marching Method (Sethian, 1996) to generate the shortest path to the goal pose. We use one of four discrete navigation actions (see Section 3.2) to reach each waypoint on the generated shortest path.
Appendix F Cross-Benchmark Experiments
In this section, we provide details about the training and evaluation setup for results on the VSI-Bench Yang et al. (2024) dataset. For evaluation, we follow the exact same evaluation protocol outlined in the original paper. In  table 3, we report results across the multiple-choice ($2490$ in number) and numerical regression ($2640$ in number) QA categories, each of which has four sub-categories. For all experiments, we use the Qwen2.5-VL-3B as the base model. Each SFT/RL-finetuning experiment uses a single node of $8$ A40 GPUs with $48$ GB VRAM each.
Supervised Finetuning. We conduct supervised finetuning of model with maximum of $48$ frames in the input context window. We also use $48$ frames during model inference as that works best between $32$, $48$ and $64$ frame inference for all baselines. For SFT experiments with the Video-R1-CoT-165k-FindingDory mixture, we experimented with various mixing ratios by changing the amount of FindingDory data ($12.5\%,25\%,37.5\%,45\%$) and found $25\%$ ($40000$ samples) to work best. All SFT models are trained for one epoch following the hyperparameters from Feng et al. (2025). We only evaluate the final checkpoint after one epoch for all baselines. Specifically, we use a learning rate of $1\times 10^{-6}$, gradient_accumulation of 2, max_grad_norm of 5 and train with bfloat16 precision on 8 GPUs We leverage gradient_checkpointing, flash_attention and deepspeed_zero_2 to optimize training.
RL Finetuning. For RL finetuning with GRPO, we reduce the context window to $32$ frames to avoid out-of-memory errors encountered with $48$ frames. Emperically, we observe that inferencing with $64$ frames works better than $32$ frame inference for all baselines. We mix the full FindingDory training dataset ($\sim 78$k samples) with the Video-R1-260k dataset from Feng et al. (2025). Because each training step is computationally expensive, we train for only $1000$ gradient steps using a group size of $8$, batch size of $1$ per GPU, and gradient accumulation of $2$. As a result, the model is exposed to only $\sim 2000$ unique samples during training. We set the weight decay to $0.01$, maximum gradient norm to $5$, learning rate to $1\times 10^{-6}$, and maximum completion length to $768$. We evaluate checkpoints at a $100$ step frequency and select the checkpoint with highest average MCQ+Numerical accuracy for each baseline.
For FindingDory samples, we design a custom reward function that rewards the precision of the predicted goal frames. Let $\mathcal{R}=(R_{1},\ldots,R_{n})$, $\mathcal{H}=(H_{1},\ldots,H_{m})$, where $R_{i}\subset\mathbb{Z}$ (reference frame sets) and $H_{i}$ are finite sequences of predicted frames.
Define the score $S(\mathcal{H},\mathcal{R})$ as
[…]
Appendix G Qualitative Examples
@@
In this section, we include representative qualitative examples from the FindingDory task suite. In Figure 16, we provide a birds-eye view depiction of the trajectory executed during experience collection (see Section B.2) in a validation episode involving 5 pick-place interaction routines. For this episode, the collected video consists of $1308$ frames in total. We provide detailed responses (and meta-analysis) from various baselines across the various task categories in Figures 18, 19, 20, 21, 22 and 23. In the following, we briefly discuss the different failure modes encountered in the generated responses across the 3 high-level task categories defined in Table 2.
@@
Single-Goal Spatial Tasks. We observe that for the Object Recall task (fig. 18), the models are able to identify a frame with the target chest of drawers but do not select the closest frames to ensure success. In the Interaction based tasks (figs. 18 and 19), we observe models fail to accurately list the sequence of interaction events which can be partially attributed to supplying only $96$ subsampled frames from the original video. In one case, GPT-4o instead focuses on a “distractor” target object (fig. 19). All models perform poorly on the Spatial Relationship tasks (fig. 19) as they are unable to coherently reason about the spatial layout of the house in the provided video.
@@
Method
All
Tasks
Object
Recall
Inter-
action
Conditional
Interaction
Object
Attributes
Spatial
Relation
Room
Visitation
Interaction
Order
Time-
Based
Duration
Tracking
Unordered
Revisitation
Ordered
Revisitation
Oracle 95.92 100.00 98.67 92.03 99.80 85.42 80.00 99.86 99.50 100.00 99.83 99.00
Text Agent 9.45 9.50 20.00 7.99 6.84 4.17 19.36 7.37 24.50 15.46 0.83 0.67
Gemma3 13.04 19.50 9.00 13.22 17.82 5.83 30.61 12.31 28.50 21.13 0.33 2.00
Qwen2.5-VL 15.14 15.50 35.00 15.22 17.62 4.58 31.83 9.83 40.00 12.76 1.83 1.33
GLM-4-1v 23.49 29.00 49.17 24.59 30.35 7.08 36.34 21.05 50.50 16.11 3.33 3.00
Gemini-2.0-Flash 25.73 47.00 33.83 30.16 35.27 10.42 41.01 23.23 43.50 21.48 8.50 9.33
GPT-4o 27.33 32.00 40.00 28.19 25.20 14.58 44.68 28.56 63.00 24.83 6.33 7.00
Qwen SFT 52.44 64.50 70.63 46.14 42.65 24.58 53.47 68.22 88.50 53.53 29.70 20.00
Table 4: High-level goal success rate (%) across task categories. The Oracle row is shaded in gray.
@@
Figure 18: FindingDory Spatial Tasks for the sample episode in Figure 16. Each VLM model uses $96$ subsampled frames. Text Agent processes full video as defined in section D.4.
Figure 19: FindingDory Spatial Tasks for the sample episode in Figure 16. Each VLM model uses $96$ subsampled frames. Text Agent processes full video as defined in section D.4.
Figure 20: FindingDory Spatial Tasks for the sample episode in Figure 16. Each VLM model uses $96$ subsampled frames. Text Agent processes full video as defined in section D.4.
Figure 21: FindingDory Temporal Tasks for the sample episode in Figure 16. Each VLM model uses $96$ subsampled frames. Text Agent processes full video as defined in section D.4.
Figure 22: FindingDory Temporal Tasks for the sample episode in Figure 16. Each VLM model uses $96$ subsampled frames. Text Agent processes full video as defined in section D.4.
Figure 23: FindingDory Multi-Goal Tasks for the sample episode in Figure 16. Each VLM model uses $96$ subsampled frames. Text Agent processes full video as defined in section D.4.